OpenStudy (anonymous):

lim x.sin3x/1-cos2x x->0

myininaya (myininaya):

Hey! I'm here

myininaya (myininaya):

\[\lim_{x->0} \frac{\sin(3x)}{1-\cos(2x)}\]

myininaya (myininaya):

Ok we will probably be using either or both of these limits: \[\lim_{u->0} \frac{\sin(u)}{u}=1\] OR \[\lim_{u->0} \frac{\cos(u)-1}{u}=0\]

myininaya (myininaya):

I think it would cool if there was someway to write the bottom as one term but we won't be able to do it without one factor Do you know what factor we need?

myininaya (myininaya):

Like we need to multiply top and bottom by ----?-----

OpenStudy (anonymous):

look! \[\lim_{x \rightarrow 0} x \sin3x/2\sin2x = xsin3x/2sinxsinx = \lim x \times \sin3x/x/2sinx \times sinx/x = 3/2\]

myininaya (myininaya):

\[1-\cos(2x)=1-(\cos^2(x)-\sin^2(x))=1-\cos^2(x)+\sin^2(x)\] \[=1-(1-\sin^2(x))+\sin^2(x)\] \[2 \sin^2(x)=2 \sin(x) \sin(x)\] Yay we can do it that was we write the bottom! :)

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(3x)}{2 \sin(x) \sin(x)}\]

myininaya (myininaya):

Ok yea I see that is what you have on the bottom at one part Good Job

myininaya (myininaya):

ok so... in order to use \[\lim_{u->0} \frac{\sin(u)}{u}=1\] We need to have sin(3x)/(3x) or even x/(sin(x)) <-we actually prove this also goes to 1 on the way to proving the first so we have \[\frac{1}{2}\lim_{x \rightarrow 0}\frac{\sin(3x)}{1} \frac{1}{\sin(x)} \frac{1}{\sin(x)}\] = \[\frac{3}{2}\lim_{x \rightarrow 0}\frac{\sin(3x)}{3x} \frac{x}{\sin(x)} \frac{1}{\sin(x)}\] the first thing I did was multiply by 3x/3x Do you see that?

OpenStudy (anonymous):

yes thanks

myininaya (myininaya):

ok so we have \[\frac{3}{2}(1)(1) \cdot \lim_{x->0}\frac{1}{\sin(x)}\]

myininaya (myininaya):

But there is no way to any manipulation so that csc(x) is continuous at x=0 The limit does not exist

OpenStudy (anonymous):

no problem may you solve it? \[\lim_{x \rightarrow 0} 2x^{2}/1-cosx\]